# Measuring and Controlling Instruction (In)Stability in Language Model Dialogs

### v3

@@ 1 Introduction
This paper presents four contributions. (1) We provide a quantitative benchmark for evaluating instruction drift that does not depend on human annotation or API calls to proprietary LLMs. This reproducible benchmark enables the measurement of progress in controlling instruction drift for both open- and closed-source models (Section 3); (2) We discuss the phenomenon of attention decay and theoretically explain why it may occur (Sections 4 and 4.3); (3) We hypothesize that attention decay is the cause of instruction drift and devise a simple technique called split-softmax as a first step towards controlling it (Section 5.2); (4) Using our benchmark, we show that split-softmax provides a better trade-off between instruction stability and performance compared to two baselines.
@@ 2 Related Work
System Prompt $s_{A}$ You are very happy! Always respond with lots of joy.
System Prompt $s_{B}$ Always reply in French.
Conversation Starter $a_{1}$ What’s your take on celebrity culture?
Probe Question $p_{B}$ What do you do in London as a tourist?
Stability measure $f_{B}(\cdot)$ [Uncaptioned image]
@@ 3.2 Benchmark Dataset
Of course, no single conversation can yield statistically significant results. To assess the degree to which a chatbot is vulnerable to instruction drift, we need to average the results of many conversations. We manually curate a benchmark set of $100$ system prompts, categorized into $5$ categories: multi-choice responses, character of the agent, answer-string format pattern, memorization of certain facts, and languages the agent speaks. Each system prompt $s_{B}$ comes with its own probe question $p_{B}$ and stability measure $f_{B}(\cdot)$, expressed as a Python function. Each stability measure $f_{B}(\cdot)$ takes as input the agent LM’s response $b_{i}$ and returns a number $p$ in the range $0\leq p\leq 1$ deterministically; the larger the value of $p$, the better the system prompt is followed. Table 1 shows one such triplet of system prompt, probe question, and stability measure. We will release the full dataset as well as the conversation starters we use.
@@ 4.2 The Phenomenon of Attention Decay
As shown in Figure 4, within each turn, $\pi(t)$ remains almost constant, but there are significant decreases across turns. This observation runs our a naive hypothesis of attention decay—if the attention distributes uniformly over previous tokens, $\pi(t)$ should decay hyperbolically and be independent of number of turns.
@@
4.3 A Geometric View of Attention Decay
To shed light on attention decay in Figure 4, both the plateau within utterance and the drop across utterances, we provide a theoretical explanation in a simplified situation. Liang et al. (2022) show empirically and theoretically that the internal representation of deep neural networks usually live in a narrow cone in the high-dimensional space. Motivated by their observations, we characterize attention decay from a similar geometric perspective.
@@ 4.3 A Geometric View of Attention Decay
For the first setting, we will show that tokens generated by the model always remain in an approximately low-dimensional convex cone in Theorem A.1. In the second setting, we can characterize the expansion using spherical measure and show that randomly drawn tokens will lead to an expansion of the underlying convex cone with the growth of intrinsic dimension of token embeddings, as shown in Proposition A.3. More details in Appendix A.
@@ 5.4 Experimental Results
6 Conclusions and Future Work
@@ Acknowledgments
KL is supported by a fellowship from the Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University. DB is supported by a grant from Open Philanthropy. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. This work was partially supported by NSF grant IIS-1901030.
Appendix A Sketch of the Theory in Section 4.3
A.1 Setting One: Agent Utterances
[…]
Theorem A.1.
Theorem A.1 shows the convex cone for token embeddings remains stable during the generating process if there is no user input, which leads to the plateau within an utterance.
A.2 Setting Two: User Utterances
@@
Appendix B Proofs for Appendix A
@@
Proof of Theorem A.1.
Let $\overline{C^{\epsilon}}$ be the convex hull of $C^{\epsilon}$. The $\overline{C^{\epsilon}}$ is a convex cone containing $C^{\epsilon}$. Theorem A.1 can be proven in two steps.
@@
In Section 4, we assume that the embedding vectors have the unit norm. To verify if this is reasonable, we plot the density of the norms of vocabulary embeddings for the LLaMA2-7B-chat in Figure 7. We can observe that the norms are quite concentrated around $1$.

### v4

@@ 1 Introduction
This paper presents four contributions. (1) We provide a quantitative benchmark for evaluating instruction drift that does not depend on human annotation or API calls to proprietary LLMs. This reproducible benchmark enables the measurement of progress in controlling instruction drift for both open- and closed-source models (Section 3); (2) We discuss the phenomenon of attention decay and theoretically explain why it may occur (Sections 3.2 and 5); (3) We hypothesize that attention decay is the cause of instruction drift and devise a simple technique called split-softmax as a first step towards controlling it (Section 6.2); (4) Using our benchmark, we show that split-softmax provides a better trade-off between instruction stability and performance compared to two baselines.
@@ 3.2 Benchmark Dataset
System Prompt $s_{A}$ You are very happy! Always respond with lots of joy.
System Prompt $s_{B}$ Always reply in French.
Conversation Starter $a_{1}$ What’s your take on celebrity culture?
Probe Question $p_{A}$ Describe the most disappointing experience you had.
Probe Question $p_{B}$ What do you do in London as a tourist?
Stability measure $f_{A}(\cdot)$ [Uncaptioned image]
Stability measure $f_{B}(\cdot)$ [Uncaptioned image]
Of course, no single conversation can yield statistically significant results. To assess the degree to which a chatbot is vulnerable to instruction drift, we need to average the results of many conversations. We manually curate a benchmark set of $100$ system prompts, categorized into $5$ categories: multi-choice responses, character of the agent, answer-string format pattern, memorization of certain facts, and languages the agent speaks. Each system prompt $s_{B}$ comes with its own probe question $p_{B}$ and stability measure $f_{B}(\cdot)$. For system prompts like $s_{A}$ in  Table 1, specific probe questions are crafted to guide the model to break the instruction; for the rest, neutral ones like $s_{B}$, a generic probe question is sampled randomly from a set of them. Each stability measure is expressed as a Python function $f_{B}(\cdot)$ that takes as input the agent LM’s response $b_{i}$ and returns a number $p$ in the range $0\leq p\leq 1$ deterministically; the larger the value of $p$, the better the system prompt is followed. Table 1 shows examples of system prompt, probe question, and stability measure. We have released the full dataset at https://huggingface.co/datasets/Naomibas/llm-system-prompts-benchmark.
@@ 4.2 The Phenomenon of Attention Decay
As shown in Figure 4, within each turn, $\pi(t)$ remains almost constant, but there are significant decreases across turns. This observation runs over a naive hypothesis of attention decay—if the attention distributes uniformly over previous tokens, $\pi(t)$ should decay hyperbolically and be independent of number of turns.
@@
5 A Geometric View of Attention Decay
To shed light on attention decay in Figure 4, both the plateau within utterance and the drop across utterances, we provide a theoretical explanation in a simplified situation.It has been shown empirically and theoretically that the internal representation of deep neural networks usually live in a narrow cone in the high-dimensional space (Mimno & Thompson, 2017; Ethayarajh, 2019; Zhu et al., 2021; Liang et al., 2022). Motivated by their observations, we characterize attention decay from a similar geometric perspective.
@@ 5 A Geometric View of Attention Decay
For the first setting, we will show that tokens generated by the model always remain in an approximately low-dimensional convex cone in Theorem 5.1. In the second setting, we can characterize the expansion using spherical measure and show that randomly drawn tokens will lead to an expansion of the underlying convex cone with the growth of intrinsic dimension of token embeddings, as shown in Proposition A.2 in Appendix A.
5.1 Setting One: Agent Utterances
[…]
Theorem 5.1.
Theorem 5.1 shows the convex cone for token embeddings remains stable during the generating process if there is no user input, which leads to the plateau within an utterance.
@@ 6.4 Experimental Results
7 Conclusions, Discussions, and Future Work
@@
The instruction drift indicates that current LM systems still cannot behave coherently over long horizons, which is the basis of long-term planning, such as information seeking (Lin et al., 2023) and tool usage (Nakano et al., 2021). Fundamentally, it is caused by a mismatch between their training schemes (text continuation or single-round RLHF) and their deployment scenarios (open-ended dialog with users). The phenomenon of instruction drift suggests there is much to be done to bridge this gap toward more robust and coherent dialog systems.
@@ Acknowledgments
KL is supported by a fellowship from the Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University and Superalignment Fast Grants from OpenAI. DB is supported by a grant from Open Philanthropy. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. This work was partially supported by NSF grant IIS-1901030.
Appendix
Appendix A Sketch of the Theory for Setting Two: User Utterances ( Section 5)
@@
Appendix B Proofs for Theorems
@@
Proof of Theorem 5.1.
Let $\overline{C^{\epsilon}}$ be the convex hull of $C^{\epsilon}$. The $\overline{C^{\epsilon}}$ is a convex cone containing $C^{\epsilon}$. Theorem 5.1 can be proven in two steps.
@@
In Section 3.2, we assume that the embedding vectors have the unit norm. To verify if this is reasonable, we plot the density of the norms of vocabulary embeddings for the LLaMA2-7B-chat in Figure 7. We can observe that the norms are quite concentrated around $1$.